A box contains ten light bulbs of which four are defective. A bulb is selected from the box and tested. If it is defective, another bulb is selected and tested, until a non-defective bulb is chosen. Let the random variable X be the number of bulbs chosen. a) Write sample space for this random experiment? b) Find probability mass Function of X? c) Find CDF of X?


1
Expert's answer
2021-01-11T18:39:18-0500

a) Sample space: "S=\{1,2,3,4,5\}"

"X\geq 1" , because we must choose at least one bulb;

We can select at most 5 bulbs, because we have only 4 defective bulbs and the last bulb chosen is non-defective.

b) If we selected "n" defective bulbs, then the probability that the next one is defective is "\frac{\text{No. of defective bulbs left}}{\text{Total no. of bulbs left}}=
\frac{4-n}{10-n}" .

If we selected "n" defective bulbs, then the probability that the next one is non-defective is "\frac{\text{No. of non-defective bulbs left}}{\text{Total no. of bulbs left}}=
\frac{6}{10-n}" .

Now we can find the probability mass Function of X:


"P(X=1)=\frac{6}{10}=\frac{3}{5}"


"P(X=2)=\frac{4}{10}\cdot \frac{6}{9}=\frac{4}{15}"


"P(X=3)=\frac{4}{10}\cdot \frac{3}{9}\cdot \frac{6}{8}=\frac{1}{10}"


"P(X=4)=\frac{4}{10}\cdot \frac{3}{9}\cdot \frac{2}{8}\cdot \frac{6}{7}=\frac{1}{35}"


"P(X=5)=\frac{4}{10}\cdot\frac{3}{9}\cdot \frac{2}{8}\cdot \frac{1}{7}\cdot \frac{6}{6}=\frac{1}{210}"


c) "F(1)=P(X\leq1)=\frac{3}{5}"


"F(2)=P(X\leq 2)=P(X=1)+P(X=2)=\frac{3}{5}+\frac{4}{15}=\frac{13}{15}"


"F(3)=P(X\leq 3)=P(X=1)+P(X=2)+P(X=3)=\frac{3}{5}+\frac{4}{15}+\frac{1}{10}=\frac{29}{30}"


"F(4)=P(X\leq 4)=P(X=1)+P(X=2)+P(X=3)+P(X=4)=\frac{3}{5}+\frac{4}{15}+\frac{1}{10}+\frac{1}{35}=\frac{209}{210}"


"F(5)=P(X\leq 5)=P(X=1)+P(X=2)+P(X=3)+P(X=4)+P(X=5)=\frac{3}{5}+\frac{4}{15}+\frac{1}{10}+\frac{1}{35}+\frac{1}{210}=1"



Answers:

a) "S=\{1,2,3,4,5\}"

b) "\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 X & 1 & 2 & 3 & 4 & 5 \\ \hline
 f(x) & \tfrac{3}{5} & \tfrac{4}{15} & \tfrac{1}{10} & \tfrac{1}{35} & \tfrac{1}{210} \\
 
\end{array}"

c) "\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 X & 1 & 2 & 3 & 4 & 5 \\ \hline
 F(x) & \tfrac{3}{5} & \tfrac{13}{15} & \tfrac{29}{30} & \tfrac{209}{210} & 1 \\
 
\end{array}"